OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution. x/2+x/3=5

OpenStudy (anonymous):

Get a common detonator: (3)x/(3)(2)+(2)x/(2)(3)=5(6)/(6) simplify: 3x/6+2x/6=30/6 add: 5x/6 =30/6 multiply by 6: 5x=30 divide by 5: x=6

OpenStudy (anonymous):

so x=6 will be our answer here?